# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

### v1

@@ 3.2 Tutorial 2: Pulling on a carbon nanotube
In this tutorial, the system of interest is a small, single-walled carbon nanotube (CNT) in an empty box (Fig. 10). The CNT is strained by imposing a constant velocity on the edge atoms. To illustrate the difference between conventional and reactive force fields, this tutorial is divided into two parts: in the first part, a conventional molecular force field (called OPLS-AA [10]) is used and the bonds between the atoms of the CNT are unbreakable. In the second part, a reactive force field (called AIREBO [11]) is used, which allows chemical bonds to break under large strain.
@@
Figure 10: The carbon nanotube (CNT) simulated during Tutorial 2.
@@ 3.2.1 Unbreakable bonds
With most conventional molecular force fields, the chemical bonds between atoms are defined at the start of the simulation and remain fixed, regardless of the forces applied to the atoms. These bonds are typically modeled as springs with equilibrium distances $r_{0}$ and force constants $k_{\text{b}}$: $U_{\text{b}}=k_{\text{b}}\left(r-r_{0}\right)^{2}$. Additionally, angular and dihedral constraints are often imposed to preserve the molecular structure by maintaining the relative orientations of neighboring atoms.
@@
The chosen unit system is real (therefore distances are in Ångströms (Å), times in femtoseconds (fs), and energies in kcal/mol), the atom_style is molecular (therefore atoms are point particles that can form bonds with each other), and the boundary conditions are fixed. The boundary conditions do not matter here, as the box boundaries were placed far from the CNT. Just like in the previous tutorial, Lennard-Jones fluid, the pair style is lj/cut (i.e. a Lennard-Jones potential with cutoff) and its cutoff is set to 14 Å, which means that only the atoms closer than this distance interact through the Lennard-Jones potential.
The bond_style, angle_style, dihedral_style, and improper_style commands specify the different potentials used to constrain the relative positions of the atoms. The special_bonds command sets the weighting factors for the Lennard-Jones interactions between atoms directly connected by one bond, two bonds, and three bonds, respectively. This is done for convenience when parameterizing the force constants for bonds, angles, and so on. By excluding the non-bonded (Lennard-Jones) interactions for these pairs, those interactions do not need to be considered when determining the force constants.
@@
Note 3.10.
@@
[…]
Note 3.11.
Rather than copying the contents of the file into the input, we incorporate it using the include command. Using include allows us to conveniently reuse the parameter settings in other inputs or switch them with others. This will become more general when using type labels, which is shown in the next tutorial [12].
@@
With the three set commands, we assign unique, otherwise unused molecule IDs to atoms in those three groups. We will use this IDs later to assign different colors to these groups of atoms.
@@
Re-setting the atom IDs is necessary before using the velocity command when atoms were deleted, which is done here with the reset_atoms command. The velocity command gives initial velocities to the atoms of the middle group cnt_mid, ensuring an initial temperature of $T=300\,\text{K}$ for these atoms.
@@
The fix nve commands are applied to the atoms of cnt_top and cnt_bot, respectively, and will ensure that the positions of the atoms from these groups are recalculated at every step. The fix nvt does the same for the cnt_mid group, while also applying a Nosé-Hoover thermostat with desired temperature of 300 K [33, 34]. To restrain the motion of the atoms at the edges, let us add the following commands to unbreakable.lmp:
@@
The two setforce commands cancel the forces applied on the atoms of the two edges, respectively. The cancellation of the forces is done at every step, and along all 3 directions of space, $x$, $y$, and $z$, due to the use of 0 0 0. The two velocity commands set the initial velocities along $x$, $y$, and $z$ to 0 for the atoms of cnt_top and cnt_bot, respectively. As a consequence of these last four commands, the atoms of the edges will remain immobile during the simulation (or at least they would if no other command was applied to them).
Note 3.12.
The velocity set command imposes the velocity of a group of atoms at the start of a run but does not enforce the velocity during the entire simulation. When velocity set is used in combination with setforce 0 0 0, as is the case here, the atoms won’t feel any force during the entire simulation. According to the Newton equation, no force means no acceleration, meaning that the initial velocity will persist during the entire simulation, thus producing a constant velocity motion.
@@
Let us import the simulation data into Python, and generate a stress-strain curve. Here, the stress is defined as $F_{\text{cnt}}/A_{\text{cnt}}$, where $A_{\text{cnt}}=\pi r_{\text{cnt}}^{2}$ is the surface area of the CNT, and $r_{\text{cnt}}=5.2$ the CNT radius. The strain is defined as $(L_{\text{cnt}}-L_{\text{cnt-0}})/L_{\text{cnt-0}}$, where $L_{\text{cnt-0}}$ is the initial CNT length.
@@
Figure 13: Stress applied on the CNT during deformation, $F_{\text{cnt}}/A_{\text{cnt}}$, where $F_{\text{cnt}}$ is the force and $A_{\text{cnt}}$ the CNT surface area, as a function of the strain, $\Delta L_{\text{cnt}}=(L_{\text{cnt}}-L_{\text{cnt-0}}/L_{\text{cnt-0}})$, where $L_{\text{cnt}}$ is the CNT length and $L_{\text{cnt-0}}$ the CNT initial length, as simulated during Tutorial 2. Here, the potential is OPLS-AA, and the CNT is unbreakable.
@@ 3.2.2 Breakable bonds
[…]
Note 3.14.
@@
[…]
@@
Figure 15: a) Evolution of the total energy $E$ of the CNT with time $t$. b) Stress applied on the CNT during deformation, $F_{\text{cnt}}/A_{\text{cnt}}$, where $F_{\text{cnt}}$ is the force and $A_{\text{cnt}}$ the CNT surface area, as a function of the strain, $\Delta L_{\text{cnt}}=(L_{\text{cnt}}-L_{\text{cnt-0}}/L_{\text{cnt-0}})$, where $L_{\text{cnt}}$ is the CNT length and $L_{\text{cnt-0}}$ the CNT initial length, as simulated during Tutorial 2. Here, the potential is AIREBO, and the CNT is breakable.
@@
This “hack” works because AIREBO does not pay any attention to bonded interactions and computes the bond topology dynamically inside the pair style. Thus adding bonds of bond style zero does not add any interactions but allow the visualization of them with dump image. It is, however needed to change the special_bonds setting to disable any neighbor list exclusions as they are common for force fields with explicit bonds.
@@ 3.3 Tutorial 3: Polymer in water
The goal of this tutorial is to use LAMMPS to solvate a small hydrophilic polymer (PEG - polyethylene glycol) in a reservoir of water (Fig. 16). Once the water reservoir is properly equilibrated at the desired temperature and pressure, the polymer molecule is added and a constant stretching force is applied to both ends of the polymer. The evolution of the polymer length is measured as a function of time. The GROMOS 54A7 force field [35] is used for the PEG, the SPC/Fw model [36] is used for the water, and the long-range Coulomb interactions are solved using the PPPM solver [37]. This tutorial was inspired by a publication by Liese and coworkers, in which molecular dynamics simulations are compared with force spectroscopy experiments, see Ref.  38.
@@ 3.3.1 Preparing the water reservoir
With the unit style real, masses are in g/mol, distances in Å, time in fs, and energies in kcal/mol. With the atom_style full, each atom is a dot with a mass and a charge that can be linked by bonds, angles, dihedrals, and/or impropers. The bond_style, angle_style, and dihedral_style commands define the potentials for the bonds, angles, and dihedrals used in the simulation, here harmonic. With the pair_style named lj/cut/coul/long, atoms interact through both a Lennard-Jones (LJ) potential and Coulomb interactions. The value of $10\,\text{\AA{}}$ is the cutoff, and the ewald command defines the long-range solver for the Coulomb interactions [40]. Finally, the special_bonds command, which was already seen in tutorial 2, sets the LJ and Coulomb weighting factors for the interaction between neighboring atoms.
@@
The overlap 1.0 option of the create_atoms command ensures that no atoms are placed exactly in the same position, as this would cause the simulation to crash. The maxtry 50 asks LAMMPS to try at most 50 times to insert the molecules, which is useful in case some insertion attempts are rejected due to overlap. In some cases, depending on the system and the values of overlap and maxtry, LAMMPS may not create the desired number of molecules. Always check the number of created atoms in the log file (or in the «Output» window), where you should see:
@@
Resetting the step of the simulation to 0 using the reset_timestep command is optional. It is used here because the number of iterations performed by the minimize command is usually not a round number, since the minimization stops when one of four criteria is reached. We will use fix npt to control the temperature and pressure of the molecules with a Nosé-Hoover thermostat and barostat, respectively [33, 34, 41]. Add the following line into water.lmp:
@@
[…]
@@
Let us also extract the volume and density every 500 steps:
variable myvol equal vol
variable myoxy equal count(H2O)/3
variable NA equal 6.022e23
variable Atom equal 1e-10
variable M equal 0.018
variable rho equal \${myoxy}*\${M}/(v_myvol*\${NA}*\${Atom}^3)
@@
thermo_style custom step temp etotal v_myvol v_rho
Here, several variables are defined and used for converting the units of the density in kg/mol: The variable myoxy represents the number of atoms divided by 3, which corresponds to the number of molecules, $N_{\text{H2O}}$, and the variable myrho is the density in kg/mol:
$\rho=\dfrac{N_{\text{H2O}}}{VN_{\text{A}}},$ (2)
where $V$ is the volume in m3, $N_{\text{A}}$ the Avogadro number, and $M=0.018$ kg/mol the molar mass of water.
@@ 3.3.2 Solvating the PEG in water
[…]
@@
Note that the recenter command has no impact on the dynamics, it simply repositions the frame of reference so that any drift of the system is ignored, which can be convenient for visualizing and analyzing the system.
@@
[…]
@@ 3.3.3 Stretching the PEG molecule
dump mydmp all local 100 pull.dat index c_dphi c_prop
@@
[…]
@@
Note 3.17.
Since the trajectory dump file does not contain information about topology and elements, it is usually preferred to first write out a data file and import it directly (in the case of OVITO) or convert it to a PSF file (for VMD). This allows the topology to be loaded before adding the trajectory file to it. When using LAMMPS–GUI, this process can be automated through the «View in OVITO» or «View in VMD» options in the «Run» menu. Afterwards only the trajectory dump needs to be added.

### v2

@@ 3.2 Tutorial 2: Pulling on a carbon nanotube
In this tutorial, the system of interest is a small, single-walled carbon nanotube (CNT) in an empty box (Fig. 10). The CNT is strained by imposing a constant velocity on the edge atoms. To illustrate the difference between conventional and reactive force fields, this tutorial is divided into two parts: in the first part, a conventional molecular force field (called OPLS-AA [10]) is used and the functional form of the bonded potential ensures that the bonds between the atoms of the CNT are unbreakable. In the second part, a reactive, many-body force field (called AIREBO [11]) is used, which allows chemical bonds to break under large strain.
Figure 10: The carbon nanotube (CNT) simulated during Tutorial 2.
@@ 3.2.1 Unbreakable bonds
With most conventional molecular force fields, the chemical bonds between atoms are defined at the start of the simulation and remain fixed, regardless of the forces applied to the atoms. In this tutorial, these bonds are explicitly specified in the .data file, which is read using the read_data command (see below). Bonds are typically modeled as springs following Hooke’s law with equilibrium distances $r_{0}$, force constants $k_{\text{b}}$, and bond potential energy $U_{\text{b}}=k_{\text{b}}\left(r-r_{0}\right)^{2}$. Additionally, angular and dihedral interactions are often imposed to preserve the molecular structure by maintaining the relative orientations of neighboring atoms.
@@
The chosen unit system is real (therefore distances are in Ångströms (Å), times in femtoseconds (fs), and energies in (kcal/mol)), the atom_style is molecular (therefore atoms are point particles that can form bonds with each other), and the boundary conditions are fixed. The boundary conditions do not matter here, as the box boundaries were placed far from the CNT. Just like in the previous tutorial, Lennard-Jones fluid, the pair style is lj/cut (i.e. a Lennard-Jones potential with cutoff) and its cutoff is set to 14 Å, which means that only the atoms closer than this distance interact through the Lennard-Jones potential.
The bond_style, angle_style, dihedral_style, and improper_style commands specify the different potentials used to constrain the relative positions of the atoms. The special_bonds command sets the weighting factors for the Lennard-Jones interactions between atoms sitting one, two, or three bonds away from each other, respectively. This is done for convenience when parameterizing the force constants for bonds, angles, and so on. By excluding the non-bonded (Lennard-Jones) interactions for these pairs, those interactions do not need to be considered when determining the force constants.
@@
Note 3.16.
Bonds, angles, dihedrals, and impropers in LAMMPS are assigned types and IDs, just like atoms. The ID uniquely identifies each interaction instance, while the type determines which parameters (from the bond_coeff, angle_coeff, etc. commands) are applied. In this tutorial, these types and IDs are specified in the .data file and read by the read_data command.
Note 3.17.
@@
The pair_coeff command sets the parameters for non-bonded Lennard-Jones interactions atom type 1 to $\epsilon_{11}=0.066\,\text{kcal/mol}$ and $\sigma_{11}=3.4\,\text{\AA {}}$. The bond_coeff provides the equilibrium distance $r_{0}=1.4\,\text{\AA {}}$ and the spring constant $k_{\text{b}}=469\,\text{kcal/mol/\AA {}}^{2}$ for the harmonic potential imposed between two neighboring carbon atoms. The potential is given by $U_{\text{b}}=k_{\text{b}}(r-r_{0})^{2}$. The angle_coeff gives the equilibrium angle $\theta_{0}$ and constant for the potential between three neighboring atoms : $U_{\theta}=k_{\theta}(\theta-\theta_{0})^{2}$. The dihedral_coeff and improper_coeff define the potentials for the constraints between 4 atoms.
Note 3.18.
Rather than copying the contents of the file into the input, we incorporate it using the include command. Using include allows us to conveniently reuse the parameter settings in other inputs or switch them with others. This will become more general when using type labels [12], which is shown in the next tutorial.
@@
Note 3.19.
So far, variables have been referenced dynamically during the run using the v_ prefix, which evaluates the variable as it evolves over time. Here, a dollar sign (\$) is used to expand the variable immediately at the time the input script is read.
@@
With the three set commands, we assign unique, otherwise unused molecule IDs to atoms in those three groups. A molecule ID is an integer that groups atoms into a ‘molecule’ for bookkeeping purposes, and can be useful for tracking and post-processing. We will use these IDs later to assign different colors to these groups of atoms.
@@
Re-setting the atom IDs is necessary before using the velocity command when atoms were deleted, which is done here with the reset_atoms command. The velocity command assigns random initial velocities to the atoms of the middle group cnt_mid from a uniform distribution, ensuring an initial temperature of $T=300\,\text{K}$ for these atoms.
@@
The fix nve commands are applied to the atoms of cnt_top and cnt_bot, respectively, and will ensure that the positions of the atoms from these groups are recalculated at every step. The fix nvt does the same for the cnt_mid group, while also applying a Nosé-Hoover thermostat with desired temperature of 300 K [34, 35].
Note 3.20.
The Nosé-Hoover thermostat only controls the temperature of the atoms belonging to the specified cnt_mid group. Atoms outside this group are not affected.
To immobilize the atoms at the edges, let us add the following commands to unbreakable.lmp:
@@
The two setforce commands cancel the forces applied on the atoms of the two edges, respectively. The cancellation of the forces is done at every step, and along all 3 directions of space, $x$, $y$, and $z$, due to the use of 0 0 0. Although the force on these atoms is set to zero, the fix stores the force vector acting on the group before cancellation, which can later be extracted for analysis (see below). The two velocity commands set the initial velocities along $x$, $y$, and $z$ to 0 for the atoms of cnt_top and cnt_bot, respectively. As a consequence of these last four commands, the atoms of the edges will remain immobile during the simulation (or at least they would if no other command was applied to them).
Note 3.21.
The velocity set command adjusts the velocities of a group of atoms immediately but has no effect during the simulation. When velocity set is used in combination with setforce 0 0 0, as is the case here, the initial velocity will persist during the entire simulation, thus producing a constant velocity motion or no motion at all.
@@
Figure 13: Stress applied on the CNT during deformation, $F_{\text{cnt}}/A_{\text{cnt}}$, where $F_{\text{cnt}}$ is the force and $A_{\text{cnt}}$ the CNT surface area, as a function of the strain, $\Delta L_{\text{cnt}}=(L_{\text{cnt}}-L_{\text{cnt-0}}/L_{\text{cnt-0}})$, where $L_{\text{cnt}}$ is the CNT length and $L_{\text{cnt-0}}$ the CNT initial length, as simulated during Tutorial 2. Here, the potential is OPLS-AA, and the CNT is unbreakable. The orange line shows the raw data, and the blue line represents a time-averaged curve.
@@
Let us import the simulation data into Python, and generate a stress-strain curve. Here, the stress is defined as $F_{\text{cnt}}/A_{\text{cnt}}$, where $A_{\text{cnt}}=\pi r_{\text{cnt}}^{2}$ is the surface area of the CNT, and $r_{\text{cnt}}=5.2$ Å the CNT radius. The strain is defined as $(L_{\text{cnt}}-L_{\text{cnt-0}})/L_{\text{cnt-0}}$, where $L_{\text{cnt-0}}$ is the initial CNT length.
@@ 3.2.2 Breakable bonds
Note 3.22.
The AIREBO force field is a many-body potential, where interactions are not only between pairs of atoms, but also triples and quadruples representing angle and dihedral interactions. This means that there are different rules for the pair_coeff command: there must be only one command that covers all permutations of atom types by using two ’*’ wildcards. After the potential file follows a list of elements. These element names are used to look up the parameter sets in the potential file. There must be a list with as many elements as atom types following the filename. In our system, there is only one atom type (1), which is mapped to the element ’C’ in the pair_coeff command. Which elements are supported is determined by the contents of the potential file.
@@
Note the relatively small timestep of $0.0005$ ps (=0.5=0.5 fs) used. Reactive force fields like AIREBO usually require a smaller timestep than conventional ones. When running breakable.lmp with LAMMPS, you can see that the temperature deviates from the target temperature of $300\,\text{K}$ at the start of the equilibration, but that after a few steps, it reaches the target value.
Note 3.24.
@@
Figure 15: a) Evolution of the total energy $E$ of the CNT with time $t$. b) Stress applied on the CNT during deformation, $F_{\text{cnt}}/A_{\text{cnt}}$, where $F_{\text{cnt}}$ is the force and $A_{\text{cnt}}$ the CNT surface area, as a function of the strain, $\Delta L_{\text{cnt}}=(L_{\text{cnt}}-L_{\text{cnt-0}}/L_{\text{cnt-0}})$, where $L_{\text{cnt}}$ is the CNT length and $L_{\text{cnt-0}}$ the CNT initial length, as simulated during Tutorial 2. Here, the potential is AIREBO, and the CNT is breakable. The orange line shows the raw data, and the blue line represents a time-averaged curve.
After equilibration, let us set the velocity of the edges equal to $75\leavevmode\nobreak\ \text{m/s}$ (or $0.75\leavevmode\nobreak\ \text{\AA {}/ps}$) and run for a longer duration than previously. Add the following lines into breakable.lmp:
@@
This “hack” works because AIREBO does not pay any attention to bonded interactions and computes the bond topology dynamically inside the pair style. Thus adding bonds of bond style zero does not add any interactions but allows the visualization of them with dump image. It is required to change the special_bonds setting to disable any neighbor list exclusions as they are common for force fields with explicit bonds.
@@ 3.3 Tutorial 3: Polymer in water
The goal of this tutorial is to use LAMMPS to solvate a small hydrophilic polymer (PEG - polyethylene glycol) in a reservoir of water (Fig. 16). Once the water reservoir is properly equilibrated at the desired temperature and pressure, the polymer molecule is added and a constant stretching force is applied to both ends of the polymer. The evolution of the polymer length is measured as a function of time. The GROMOS 54A7 force field [36] is used for the PEG, the SPC/Fw model [37] is used for the water, and the long-range Coulomb interactions are solved using the PPPM solver [38]. This tutorial was inspired by a publication by Liese and coworkers, in which molecular dynamics simulations are compared with force spectroscopy experiments, see Ref.  39.
Note 3.25.
When mixing different force fields, as is done here with GROMOS and SPC/Fw, users should exercise caution. The choices made in these tutorials prioritize progressive learning of LAMMPS functionality over strict physical accuracy. While GROMOS is commonly used with water models from the SPC family [40], the inter-compatibility of force fields is not generally guaranteed.
@@ 3.3.1 Preparing the water reservoir
With the unit style real, masses are in g/mol, distances in Å, time in fs, and energies in kcal/mol. With the atom_style full, each atom is a dot with a mass and a charge that can be linked by bonds, angles, dihedrals, and/or impropers. The bond_style, angle_style, and dihedral_style commands define the potentials for the bonds, angles, and dihedrals used in the simulation, here harmonic. With the pair_style named lj/cut/coul/long, atoms interact through both a Lennard-Jones (LJ) potential and Coulomb interactions. The value of $10\,\text{\AA {}}$ is the cutoff, and the kspace_style command defines the long-range solver for the Coulomb interactions [42]. Finally, the special_bonds command, which was already seen in Tutorial 2, sets the LJ and Coulomb weighting factors for the interaction between neighboring atoms.
Note 3.26.
With Coulomb interactions, additional rules apply to the pair_coeff command: (a) atom type values only matter for assignment of LJ potential parameters; (b) for Coulomb interactions, there are no parameters outside the cutoff, and when using a coul/long pair style, that cutoff can only be set globally for all atoms with the pair_style command; (c) for Coulomb interactions, only the per-atom charge and any special_bonds exclusions are relevant.
@@
The first parameter is 0, meaning that the atom IDs from the water.mol file will be used. The overlap 1.0 option of the create_atoms command ensures that no atoms are placed exactly in the same position, as this would cause the simulation to crash. The maxtry 50 asks LAMMPS to try at most 50 times to insert the molecules, which is useful in case some insertion attempts are rejected due to overlap. In some cases, depending on the system and the values of overlap and maxtry, LAMMPS may not create the desired number of molecules. Always check the number of created atoms in the log file (or in the «Output» window), where you should see:
@@
Resetting the step of the simulation to 0 using the reset_timestep command is optional. It is used here because the number of iterations performed by the minimize command is usually not a round number, since the minimization stops when one of four criteria is reached, which can disrupt the intended frequency of outputs such as dump commands that depend on the timestep count. We will use fix npt to control the temperature and pressure of the molecules with a Nosé-Hoover thermostat and barostat, respectively [34, 35, 43]. Add the following line into water.lmp:
@@
The fix npt allows us to impose both a temperature of $300\,\text{K}$ (with a damping constant of $100\,\text{fs}$), and a pressure of 1 atmosphere (with a damping constant of $1000\,\text{fs}$). With the iso keyword, the three dimensions of the box will be re-scaled isotropically, maintaining the same proportion in all directions.
@@
Let us also extract the volume and density, among others, every 500 steps:
@@
thermo_style custom step temp etotal vol density
With the real units system, the volume is in Å3, and the density is in g/cm3.
@@ 3.3.2 Solvating the PEG in water
Let us also use the recenter command to always keep the PEG at the position (0, 0, 0)($\leavevmode\nobreak\ 0,\leavevmode\nobreak\$0):
@@
Note 3.29.
Note that the recenter command has no impact on the dynamics, it simply repositions the frame of reference so that any drift of the system is ignored, which can be convenient for visualizing and analyzing the system. However, be aware that using fix recenter can sometimes mask underlying issues in the simulation, such as a net momentum or the so-called “flying ice cube syndrome” [7].
@@
Run the simulation using LAMMPS. From the outputs, you can make sure that the temperature remains close to the target value of $300\leavevmode\nobreak\ \text{K}$ throughout the entire simulation, and that the volume and total energy are almost constant, indicating that the system was in a reasonable configuration from the start. See a snapshot of the system in Fig. 20.
@@ 3.3.3 Stretching the PEG molecule
dump mydmp all local 100 pull.dat index c_dphi
@@
Each applied force has a magnitude of $10\,\text{kcal/mol/\AA {}}$, corresponding to $0.67\,\text{nN}$. This value was chosen to be sufficiently large to overcome both the thermal agitation and the entropic contributions from the molecules.
@@
Note 3.30.
Since the default trajectory dump file does not contain information about topology and elements, it is usually preferred to first write out a data file and import it directly (in the case of OVITO) or convert it to a PSF file (for VMD). This allows the topology to be loaded before adding the trajectory file to it. When using LAMMPS–GUI, this process can be automated through the «View in OVITO» or «View in VMD» options in the «Run» menu. Afterwards only the trajectory dump needs to be added. Alternatively, the dump custom command can be combined with dump command to include element names in the dump file and simplify visualization.
Note 3.31.
Microstates collected during a simulation in the form of a trajectory can be analyzed within LAMMPS using the rerun command. This is particularly useful, for example, for computing properties not set up in the original simulation without having to run it again. A possible use of the rerun command is estimating the self-diffusion coefficient by using the compute msd command [1].